### v2

# The equational theory of the Weihrauch lattice with multiplication

Abstract
We study the equational theory of the Weihrauch lattice with multiplication, meaning the collection of equations between terms built from variables, the lattice operations $\sqcup$,$\sqcap$ , ⊓, the product $\times$, and the finite parallelization which are true however we substitute Weihrauch degrees for the variables. We provide a combinatorial description of these in terms of a reducibility between finite graphs, and moreover, show that deciding which equations are true in this sense is complete for the third level of the polynomial hierarchy.
Erratum
The first version of this preprint claimed that our theory was not complete because $a\times(b\sqcap c)\leq a\times(b\sqcap(c\times a))$ was not derivable without the pointedness axiom. This is false. We have removed the erroneous statement and added a proof that the axiomatization we propose is complete for Weihrauch degrees if and only if it is complete for pointed degrees if $1\leq a$ is added (Theorem 15).
1 Introduction
The Weihrauch degrees $\mathfrak{W}$ come with a rich algebraic structure. Here, we consider the lattice operations $\sqcup$,$\sqcap$ , ⊓, the product $\times$ and the finite parallelization . They were the first operations on Weihrauch degrees studied in the literature [3, 2, 14]. In order to better understand the structure of the Weihrauch degrees, we would like to characterize its equational theory, i.e. we want to identify which equations between terms over the signature $(\mathfrak{W},\sqcap,\sqcup,\times,1,(-)^{*})$ are true for every instantiation of the variables by Weihrauch degrees. It was already observed in [5] that the equational theory of $(\mathfrak{W},\sqcap,\sqcup)$ is the theory of distributive lattices, as $(\mathfrak{W},\sqcap,\sqcup)$ is a distributive lattice itself, and every countable distributive lattice embeds into $(\mathfrak{W},\sqcap,\sqcup)$ (via the Medvedev degrees).
Context
Shortly after the precise definition of Weihrauch reducibility was proposed in [9], Brattka and Gherardi raised the question whether the Weihrauch degrees $\mathfrak{W}$ form a Heyting or Brouwer algebra, in line with the overall idea that the structure of the Weihrauch degrees should be some kind of structure of constructive truth values. The question was answered in the negative in [12]; in particular, the Weihrauch degrees are not complete.
The broader question to what extend the Weihrauch degrees form an instance of already studied classes of structures linked to (constructive) logics remains an active area of research. While similarities and tentative connections are easy to find, a satisfactory answer is still eluding the community. A better understanding of the structure of the Weihrauch degrees seems integral to further progress, in particular for the hope to find the Weihrauch degrees to even be universal for some such logic.
Our contributions
We start by investigating the equational theory of $(\mathfrak{W},\sqcap,\times)$. While we only conjecture that the axiomatization we propose for it is complete, we provide a combinatorial characterization in terms of reductions between finite graphs. This in turn will allow us to show that determining universal validity of equations in $(\mathfrak{W},\sqcap,\times)$ is $\Sigma^{p}_{2}$-complete. Since we only work with substructures that include $\sqcap$, we can write inequalities instead of equations by treating $t\leq u$ as an abbreviation for $t=t\sqcap u$.
A Weihrauch degree is called pointed if its representatives have some computable instance. We denote by $\mathfrak{W}_{\bullet}\subset\mathfrak{W}$ the subset of pointed degrees and write $\leq_{\bullet}$ and $=_{\bullet}$ instead of $\leq$ and $=$ when we mean inequalities and equations that are only meant to be valid for pointed degrees. One easy observation is that if an (in)equality holds in $\mathfrak{W}$, it also holds in $\mathfrak{W}_{\bullet}$.
Occasionally a variant of Weihrauch reducibility called strong Weihrauch reducibility is studied. Our operations are also operations on strong Weihrauch degrees, which we shall denote by $\mathfrak{W}^{s}$, and it turns out that the equational theory of $(\mathfrak{W}^{s},\sqcap,\times)$ and $(\mathfrak{W},\sqcap,\times)$ coincide. However, as the strong Weihrauch degrees are not distributive as a lattice [8], this does not extend to signatures including $\sqcap$ and $\sqcup$.
Other work on the structure of the Weihrauch degrees
Another investigation into the structure of the Weihrauch degrees is [13]. A key result there is that $1$ is definable in just $(\mathfrak{W},\leq_{\textrm{W}})$ – but this involves multiple quantifiers, and thus is not directly relevant to the equational theory we study here.
At first glance, [11] seems to have a similar approach to our work. Hertling and Selivanov consider the complexity of a combinatorial reducibility which earlier work by Hertling [10] tied to continuous Weihrauch reducibility. However, nothing seems to be known about definability of the fragment of the continuous Weihrauch degrees which is characterized by the combinatorial reducibility.
Related work in proof theory
We are not aware of axioms systems matching exactly those we provided for $(\mathfrak{W}_{\bullet},\sqcap,\times)$ in the literature. Despite of this, there are tentative connections to be made between notions of correctness in substructural logics being phrased in terms of graphs and our notion of reduction in terms of graphs. In particular, [6] has a strikingly similar notion of reducibility between graphs generalizing boolean formulas. this seems to indicate casting our axioms or our general notion of graph reducibility in a deep inference system might be of relevance in uncovering a complete axiomatization for the theories we consider.
2 Introducing the operations
A general reference on Weihrauch degrees is the survey article [4]. Here, we will recap briefly the operations we study. As we are interested in only the structure of the Weihrauch degrees, we can avoid introducing represented spaces, and instead work with the following:
Definition 1.
Let $f,g:\subseteq{\{0,1\}^{\mathbb{N}}}\rightrightarrows{\{0,1\}^{\mathbb{N}}}$ be multivalued functions. We write $f\leq_{\textrm{W}}g$ and say that $f$ is Weihrauch reducible to $g$ if there are computable $H:\operatorname{dom}(f)\to\operatorname{dom}(g)$ and (partial) $K:{\{0,1\}^{\mathbb{N}}}\times{\{0,1\}^{\mathbb{N}}}\rightharpoonup{\{0,1\}^{\mathbb{N}}}$ such that for all $x\in\operatorname{dom}(f)$ and $y\in g(H(x))$ we have that $K(x,y)$ is defined and $K(x,y)\in f(x)$.
The Weihrauch degrees are the equivalence classes for $\leq_{\textrm{W}}$. We denote them by $\mathfrak{W}$. If $K$ does not depend on its first input, we have a strong Weihrauch reduction $f\leq_{\textrm{sW}}g$. The strong Weihrauch degrees are denoted by $\mathfrak{W}^{s}$.
Let :$\langle\ \rangle:{\{0,1\}^{\mathbb{N}}}\times{\{0,1\}^{\mathbb{N}}}\to{\{0,1\}^{\mathbb{N}}}$ be a standard pairing function. The operations we investigate are the following:
Definition 2.
Given multivalued functions $f,g:\subseteq{\{0,1\}^{\mathbb{N}}}\rightrightarrows{\{0,1\}^{\mathbb{N}}}$ , we define
  • $f\sqcap g$ by $0p\in(f\sqcap g)(\langle q_{0},q_{1}\rangle)$ if $p\in f(q_{0})$ and $1p\in(f\sqcap g)(\langle q_{0},q_{1}\rangle)$ if $p\in g(q_{1})$;
  • $f\sqcup g$ by $0p\in(f\sqcup g)(0q)$ if $p\in f(q)$ and $1p\in(f\sqcup g)(1q)$ if $p\in g(q)$;
  • $f\times g$ by $\langle p_{0},p_{1}\rangle\in(f\times g)(\langle q_{0},q_{1}\rangle$ if $p_{0}\in f(q_{0})$ and $p_{1}\in g(q_{1})$.
These operations on multivalued functions induce operations on Weihrauch degrees. Moreover, the operations are monotone. The Weihrauch degrees are a lattice, with $\sqcap$ as meet and $\sqcup$ as join.
We also make use of the constant $1$, which is the Weihrauch degree of $\operatorname{id}:{\{0,1\}^{\mathbb{N}}}\to{\{0,1\}^{\mathbb{N}}}$. The representatives of $1$ are exactly the computable multivalued functions whose domain contains a computable point. The upper cone of $1$ (i.e. $\{f\mid 1\leq_{\textrm{W}}f\}$) is the class of pointed Weihrauch degrees $\mathfrak{W}_{\bullet}$.
Definition 3.
Given a multivalued function $f:\subseteq{\{0,1\}^{\mathbb{N}}}\rightrightarrows{\{0,1\}^{\mathbb{N}}}$, we define $f^{*}$ by $0^{\omega}\in f^{*}(1p)$ and $\langle p_{0},p_{1}\rangle\in f^{*}(0^{n+1}1\langle q_{0},q_{1}\rangle)$ if $p_{0}\in f(q_{0})$ and $p_{1}\in f^{*}(0^{n}1q_{1})$.
Plainly spoken, $f^{*}$ receives a number $n\in\mathbb{N}$ together with $n$ inputs for $f$ and outputs $n$ corresponding solutions.
Basic properties of $(\mathfrak{W},\sqcap,\times,1)$
{mdframed}
$\begin{array}[]{cr}a\times(b\times c)=(a\times b)\times c\qquad a\times 1=a&\text{monoid structure}\\ a\times b=b\times a&\text{commutativity}\\ a\leq a\times a&\text{relevance}\\ a\sqcap b\leq a\qquad a\sqcap b\leq b&\text{$\sqcap$ is a lower bound}\\ a\leq b\;\wedge\;a\leq c~{}\Rightarrow~{}a\leq b\sqcap c&\text{$\sqcap$ is thegreatest lb}\\ (a\times b)\sqcap c\leq a\times(b\sqcap c)&\text{half-distributivity}\end{array}$
Additional axiom for the pointed case $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$mdframed}
$\begin{array}[]{cr}1\leq_{\bullet}a&\text{bottom element}\\ \end{array}$
Figure 1: Our set of axioms for the systems under consideration, where we implicitly assume that equality is an equivalence relation, that all operators are monotone in each of their arguments and that inequalities are transitive. Every axiom we state holds when substituting $\leq_{\bullet}$ for $\leq$ everywhere, and we take this as axioms for $\leq_{\bullet}$ alongside the specific $\leq_{\bullet}$ axioms.
At first, we investigate the structures $(\mathfrak{W},\sqcap,\times,1)$ and $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$. A partial axiomatization for these is provided in Figure 1. We do not explicitly list monotonicity of the operations and transitivity of $\leq$. To express the axioms, we will only need Horn clauses, that is formulas of the shape $\bigwedge_{i}t_{i}\leq u_{i}\Rightarrow t^{\prime}\leq u^{\prime}$. Our partial axiomatization essentially states that $\sqcap$ does behave like a greatest lower bound, that $\times$ and $1$ equip $\mathfrak{W}$ with a relevant commutative ordered monoid structure and that $\times$ half-distributes over $\sqcap$. Proofs that they hold in $\mathfrak{W}$ are not difficult and can be found in [5].
We were unable to find a true inequality in $(\mathfrak{W},\sqcap,\times,1)$ that does not follow from the stated axioms, and thus offer up the following:
Conjecture 4.
$(\mathfrak{W},\sqcap,\times,1)$ is fully axiomatized by Figure 1.
We also show later that the axiomatization is complete for $\mathfrak{W}$ if and only if it the pointed version is complete for $\mathfrak{W}_{\bullet}$ (Theorem 15).
3 A combinatorial notion of reducibility for terms
The goal of this section is to give a combinatorial criterion to check the validity of an inequality. Along the way we will see that the validity of an inequality where the semantics of the variables may range over all possible degrees can be reduced to the validity of the inequality where the variables are interpreted as sufficiently independent degrees. Let us first formally state the notion we investigate:
Definition 5.
Let $u$ and $v$ be terms constructed from variables, the constant $1$ and the binary connectives ,$\sqcap$,$\times$ , ×. We say that an inequality $t\leq u$ (respectively $t\leq_{\bullet}u$) is universally valid when for every interpretation of the variable as Weihrauch degrees (respectively pointed Weihrauch degrees), there is a corresponding Weihrauch reduction.
Clearly, if $t\leq u$ is universally valid, then so is $t\leq_{\bullet}u$. Conversely, if $t\leq_{\bullet}u$ is universally valid, then $t\leq u$ is universally valid if and only if every variable occurring in $u$ also occurs in $t$. The reason for this is that the only difference between a Weihrauch reduction to $f$ and a Weihrauch reduction to $f\sqcup 1$ is that for the latter we do not necessarily need access to an instance for $f$. If every variable in $u$ appears also in $t$, then for every instantiation of the variables, we have available to us valid inputs for all Weihrauch degrees.
Now we prepare to define a combinatorial reduction between terms. To do so, we first need to introduce an interpretation of terms as graphs.
Definition 6.
A (finite coloured undirected) graph is a triple $(V,E,c)$ where $V$ is a finite set of vertices, $E\subseteq[V]^{2}$ is a set of edges and $c:V\to\mathcal{V}$ a colouring of the vertices by variables.
Definition 7.
Write $G+H$ for the disjoint union of two graphs and $G^{\bot}$ for $(V,[V]^{2}\setminus E,c)$ when $G=(V,E,c)$. The interpretation $\llbracket t\rrbracket$ of terms $t$ as graphs is defined by induction as follows:
  • for a variable $v$, $\llbracket v\rrbracket$ is a graph with a single vertex with colour $v$
  • $\llbracket t\sqcap u\rrbracket=\llbracket t\rrbracket+\llbracket u\rrbracket$ and $\llbracket t\times u\rrbracket=(\llbracket t\rrbracket^{\bot}+\llbracket u\rrbracket^{\bot})^{\bot}$
The graphs that arise as $\llbracket u\rrbracket$ for some term $u$ are known as cographs. They can also be characterized as those graphs that do not have an induced subgraph forming a line of length $3$ [7]. Below we give some examples of terms and corresponding cographs.
The intuition is that a coloured graph $(V,E,c)$ can be read as the following Weihrauch problem, assuming the colours correspond themselves to Weihrauch problems: given for each vertex $v\in V$ an input $i_{v}$ to $c(v)$, output a maximal clique $C$ of $(V,E)$ together with solutions for every $i_{v}$s with $v\in C$. Under this interpretation, the problems denoted $\llbracket t\rrbracket$ and $t$ are Weihrauch-equivalent. This leads naturally to the following notion of combinatorial reduction between coloured graphs.
Definition 8.
Given two graphs $G=(V_{G},E_{g},c_{G})$ and $H=(V_{H},E_{H},c_{H})$, call a partial function $f:V_{H}\rightharpoonup V_{G}$ a $\bullet$-reduction of $G$ to $H$ if it respects colours (i.e., $c_{H}(v)=c_{G}(f(v))$ for all $v\in V_{H}$) and the image of every maximal clique in $H$ under $f$ contains a maximal clique of $G$. If $f$ is total, it is a reduction.
An inequality $t\leq u$ (respectively $t\leq_{\bullet}u$) is said to be combinatorially valid if there is a reduction of $\llbracket t\rrbracket$ into $\llbracket u\rrbracket$ (respectively a $\bullet$-reduction).
Let us illustrate the notion of combinatorial reduction on our examples.
In the two examples above, we have a strict reduction. We do not specify the total maps between vertices as they are uniquely determined by the colors in these particular cases. The shaded areas correspond to maximal cliques in the graphs and they have matching colors when the reductions relate them. There are no reductions the other ways around: for the first scenario, this is because the clique $\{a,c\}$ on the left can only be mapped to a strict subclique of the triangular clique on the right by a color-preserving map. For the second case, where we have a unique color preserving-map, it is because $\{a,b\}$ does not contain a maximal clique in the rightmost graph.
We are now ready to state the main theorem of this section:
Theorem 9.
Let $t$ and $u$ be terms constructed from variables and the binary connectives ,$\sqcap$,$\times$ , ×. Then inequalities $t\leq u$ and $t\leq_{\bullet}u$ are universally valid if and only if they are combinatorially valid.
To prove the theorem, we shall introduce a third notion of validity. Assume that we have a countable enumeration $(x_{n})_{n\in\mathbb{N}}\in\mathcal{V}^{\mathbb{N}}$ of all the variables that we may ever want to use in our terms. We pick a countable strong antichain $(a_{n,m})_{(n,m)\in\mathbb{N}^{2}}$ of elements of ${\{0,1\}^{\mathbb{N}}}$ in the Turing degrees.
Definition 10.
We say that an inequality $t\leq u$ is generically valid if it holds when we instantiate the $n$-th variable by the function $(a_{n,0},k)\mapsto a_{n,k+1}:\{a_{n,0}\}\times\mathbb{N}\to{\{0,1\}^{\mathbb{N}}}$. We say that $u\leq_{\bullet}v$ is generically valid if it holds when we instantiate the $n$-th variable by the function $k\mapsto a_{n,k}:\mathbb{N}\to{\{0,1\}^{\mathbb{N}}}$.
Lemma 11.
A generically valid inequality $t\leq u$ or $t\leq_{\bullet}u$ is also combinatorially valid.
Proof.
That $t\leq u$ or $t\leq_{\bullet}u$ respectively is generically valid is stating that there exists a Weihrauch reduction $f_{t}\leq_{\textrm{W}}g_{u}$ between particular multivalued functions constructed from the terms. We fix an input to $f_{t}$ where different occurrences of the same variable receive different integer inputs. The valid solutions for $f$ are certain combinations of values $a_{i,j}$ from our chosen strong antichain. Following the construction of $\llbracket t\rrbracket$, we see that the sufficient combinations of $a_{i,j}$ correspond to the maximal cliques in $\llbracket t\rrbracket$.
Since the reduction cannot obtain one of the $a_{i,j}$ unless the value is returned by $g_{u}$, we can obtain the reduction from $\llbracket t\rrbracket$ to $\llbracket u\rrbracket$ by observing which values provided by $g_{u}$ are the ones demanded by $f_{t}$.
If we are considering $t\leq_{\bullet}u$, we are done. For $t\leq u$ we observe that the presence of $a_{n,0}$ in the domains means that $\llbracket u\rrbracket$ cannot have any colours which are absent from $\llbracket t\rrbracket$. We can thus make the partial reduction total by assigning some arbitrary vertex of the correct colour for any undefined output. ∎
Lemma 12.
A combinatorially valid inequality $t\leq u$ or $t\leq_{\bullet}u$ is also universally valid.
Proof.
The forward reduction is entirely determined by the combinatorial one, and the backwards strong reduction can be obtained by taking a choice function between maximal cliques corresponding to the correctness criterion for the combinatorial reduction. ∎
Proof of LABEL:{lem:combred-valid}.
It is relatively straightforward to prove
[…]
$(1)$ is trivial. $(2)$ is the statement of Lemma 11. $(3)$ is the statement of Lemma 12. ∎
Extending the picture to strong Weihrauch reducibility
The proof of Lemma 12 proceeded by observing that a combinatorial reduction between terms translates to a Weihrauch reduction between the Weihrauch degrees obtained by substituting some arbitrary multivalued functions for the variables. The Weihrauch reductions we get are even strong Weihrauch reductions. Thus, the results of LABEL:{lem:combred-valid} still hold when we replace universally valid by its counterpart for strong Weihrauch reducibility. This yields:
Corollary 13.
$(\mathfrak{W},\sqcap,\times)$ and $(\mathfrak{W}^{s},\sqcap,\times)$ have the same equational theory.
On the constant $1$
Extending our results for inequalities $t\leq_{\bullet}u$ built from ,$\sqcap$,$\times$ , × to terms also including the constant $1$ is straightforward, as our axioms for the pointed degrees allow us to simplify any term over (,×$(\sqcap,\times,1)$ to produce either just $1$ or a term over (,×)square-intersection($\sqcap$,$\times$)( ⊓ , × ).
For inequalities $t\leq u$, this is less straightforward. We can extend our interpretations of terms as coloured cographs to include $1$ by mapping $1$ to an uncoloured vertex. In the definition of a reduction map the uncoloured vertices are ignored other than in defining what the maximal cliques are. This means that the reduction map does not have to be defined on the uncoloured vertices, and the image of a maximal clique in $\llbracket u\rrbracket$ has to contain – up to some uncoloured vertices – a maximal clique in $\llbracket t\rrbracket$.
While having $1$ available to us makes it trivial to reduce universal validity of terms $t\leq_{\bullet}u$ to universal validity of terms $t^{\prime}\leq u^{\prime}$, this is possible even without $1$:
Proposition 14.
Let $t,u$ be terms built using the variables $x_{0},\ldots,x_{n}$. The term $t\leq_{\bullet}u$ is universally valid iff $t\times x_{0}\times\ldots\times x_{n}\leq u\times x_{0}\times\ldots\times x_{n}$ is universally valid.
Proof.
Write $(V_{t},E_{t},c_{t})$ and $(V_{u},E_{u},c_{u})$ for $\llbracket t\rrbracket$ and $\llbracket u\rrbracket$ respectively, and assume that we have that the vertices of $\llbracket x_{0}\times\ldots\times x_{n}\rrbracket$ being $\{0,\ldots,n\}$ and disjoint from $V_{t}$ and $V_{u}$. For both directions, we use the combinatorial validity of the reductions, which we know to be equivalent to universal validity.
If we have a pointed combinatorial reduction $t\leq_{\bullet}u$ witnessed by a partial map $f:V_{u}\rightharpoonup V_{t}$, then it can be extended to a non-pointed one witnessed by $\hat{f}:V_{u}\cup\{0,\ldots,n\}\to V_{t}\cup\{0,\ldots,n\}$ extending $f$ by setting $\hat{f}(i)=i$ for $0\leq i\leq n$ and $\hat{f}(v)=i$ when $v\in V_{u}\setminus\operatorname{dom}(f)$ and $c_{t}(v)=x_{i}$. It is easily checked to map maximal cliques to sets containing maximal cliques.
Conversely, if we have a combinatorial reduction $t\times x_{0}\times\ldots\times x_{n}\leq u\times x_{0}\times\ldots\times x_{n}$ witnessed by $g:V_{u}\cup\{0,\ldots,n\}\to V_{t}\cup\{0,\ldots,n\}$, we may massage it into a partial map $\hat{g}:V_{u}\rightharpoonup V_{t}$ witnessing a combinatorial reduction $t\leq_{\bullet}u$. $\hat{g}$ is defined as follows for $v\in V_{u}$:
  • If $g(v)\in V_{t}$, we take $\hat{g}(v)=g(v)$
  • If $g(v)=i\in\{0,\ldots,n\}$ and $g(i)\in V_{t}$, we take $\hat{g}(v)=g(i)$
  • Otherwise $v\notin\operatorname{dom}(\hat{g})$
If we have a maximal clique $M$ in $\llbracket u\rrbracket$, then $M\cup\{0,\ldots,n\}$ is a maximal clique in $\llbracket u\times x_{0}\times\ldots\times x_{n}\rrbracket$ and thus gets mapped by $g$ to a set $S$ containing a maximal clique in $\llbracket t\times x_{0}\times\ldots\times x_{n}\rrbracket$ by $g$. Writing $S^{\prime}$ for $S\cap V_{t}$, we clearly have that $S=S^{\prime}\cup\{0,\ldots,n\}$ and that $S^{\prime}$ contains a maximal clique of $\llbracket t\rrbracket$. But then we also have $\hat{g}$ maps $M$ to $S^{\prime}$: let us prove this by contradiction. If an element $e$ of $S^{\prime}$ is not in $\hat{g}(M)$, it means that there is no $v\in V_{u}$ such that $g(v)=e$ but there is some $i\in\{0,\ldots,n\}$ such that $g(i)=e$ by the first clause in the definition of $\hat{g}$. But since $i\in S$, there must be some $v^{\prime}$ in $M$ such that $g(v^{\prime})=i$. But then that would mean that $\hat{g}(v^{\prime})=e$ by the second clause, a contradiction. ∎
4 Completeness does not depend on pointedness
Theorem 15.
The axioms of Figure 1 are complete for $(\mathfrak{W},\sqcap,\times,1)$ if and only if they are complete for $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$ when adding $1\leq_{\bullet}a$
To prove this theorem, we first establish some results for derivability that allows us to relate the pointed and non-pointed theories. For any finite set $X$ of variables, call $i_{X}$ the formula $1\sqcap\bigsqcap_{x\in X}x$. The intuition is that the corresponding Weihrauch problem requires at least an input for every $x$ to be provided, but does not answer any of the corresponding questions.
Lemma 16.
For any finite sets of variables $X$ and $Y$, we can derive $i_{X}\times i_{Y}=i_{X\cup Y}$.
Proof.
Firstly we can show that we have $i_{X}\times i_{X}=i_{X}$. This is because we have the relevance axiom and $i_{X}\times i_{X}\leq i_{X}\times 1=i_{X}$, since $i_{X}\leq 1$ by definition. Then, note that $i_{X^{\prime}}\leq i_{X}$ whenever $X\subseteq X^{\prime}$. So we have $i_{X\cup Y}=i_{X\cup Y}\times i_{X\cup Y}\leq i_{X}\times i_{Y}$. For the converse direction, it suffices to show that $i_{X}\times i_{Y}\leq 1$, $i_{X}\times i_{Y}\leq x$ for all $x\in X$ and, symmetrically, $i_{X}\times i_{Y}\leq y$ for all $y\in Y$. This is straightforward as $i_{X}\leq 1$ and $i_{X}\leq x$ for every $x\in X$. ∎
Lemma 17.
For any term $t$ whose set of free variables is $X$, we can derive $i_{X}\times t=t$.
Proof.
Deriving $t\leq i_{X}\times t$ is easy since $1\leq i_{X}$. For the converse, we proceed by induction over $t$. If $t=1$, this amounts to $1\times 1=1$. If $t=t_{1}\times t_{2}$, call $X_{1}$ and $X_{2}$ the free variables of $t_{1}$ and $t_{2}$ respectively. In this case we have $X=X_{1}\cup X_{2}$, so we can derive $i_{X}\times t\leq i_{X_{1}}\times t_{1}\times i_{X_{2}}\times t_{2}$ using Lemma 16 and then conclude using the inductive hypotheses and the monotonicity axiom for $\times$. If we have $t=t_{1}\sqcap t_{2}$, adopting similar naming conventions, we also have $i_{X}\times t\leq(i_{X_{1}}\times t_{1})\sqcap(i_{X_{2}}\times t_{2})$ using the fact that $i_{X}\leq i_{Y}$ for any $Y\subseteq X$, and we similarly using monotonicity of $\sqcap$. ∎
Lemma 18.
For any terms $t$ and $u$ such that all the free variables of $t$ and $u$ are contained in a finite set $X$, then we have that $i_{X}\times t\leq u$ is derivable if and only if $t\leq_{\bullet}u$ is derivable.
Proof.
The left-to-right direction is straightforward, as we have $1\leq_{\bullet}x$ for all $x\in X$ and therefore $1\leq_{\bullet}i_{X}$. For the right-to-left direction, we need to reason by induction on the size of a derivation $t\leq_{\bullet}u$ and make a case analysis on the last axiom applied.
  • If the last axiom applied is reflexivity and $t=u$, we can simply apply Lemma 17 since $X$ is a superset of the free variables of $t$.
  • If the last axiom applied is transitivity, i.e. we have two subderivations of $t\leq_{\bullet}v$ and $v\leq_{\bullet}u$ for some $v$, we can assume, without loss of generality, that all variables occurring in $v$ occur in $X$. Because if we had such a variable $a$ in $v$, we could get derivations of the same size of $t=t[1/a]\leq_{\bullet}v[1/a]$ and $v[1/a]\leq_{\bullet}u[1/a]=u$. Therefore, we can apply the induction hypothesis to get derivations of $i_{X}\times u\leq v$ and $i_{X}\times v\leq u$ which we can chain as follows
    $i_{X}\times t\leq i_{X}\times(i_{X}\times t)\leq i_{X}\times v\leq u$
  • If the last axiom applied is monotonicity of of $\times$, i.e. we have derivations $t_{1}\leq_{\bullet}u_{1}$, $t_{2}\leq_{\bullet}u_{2}$ and $t=t_{1}\times t_{2}$ and $u=u_{1}\times u_{2}$, then we can apply the inductive hypotheses to obtain $i_{X}\times t_{j}\leq u_{j}$ for $j\in\{1,2\}$ and conclude by
    $i_{X}\times t_{1}\times t_{2}\leq i_{X}\times t_{1}\times i_{X}\times t_{2}\leq u_{1}\times u_{2}$
  • Monotonicity of $\sqcap$ is handled similarly.
  • All of the other axioms shared between the pointed and non-pointed cases are handled trivially. So only $1\leq_{\bullet}t$ remains, for which it suffices to prove that $i_{X}\leq t$. This is done via an easy induction over $t$.
Proof of Theorem 15.
First let us assume that our theory is complete for $\mathfrak{W}_{\bullet}$ and that we have that $t\leq u$ combinatorially valid in $\mathfrak{W}$. In particular, our theory derives $t\leq_{\bullet}u$. Via combinatorial validity, we also know that means that the free variables of $u$ are also free variables of $t$; call $X$ the set of free variables of $t$. By Lemma 18, we have that our theory derives $i_{X}\times t\leq u$, but we also have by Lemma 17 that $t\leq i_{X}\times t$, so we can conclude that the theory is also complete for $\mathfrak{W}$.
Conversely, if the theory is complete for $\mathfrak{W}$ and $t\leq_{\bullet}u$ is combinatorially valid in $\mathfrak{W}_{\bullet}$, then, if $X$ is the set of free variables of $u$, $i_{X}\times t\leq u$ is easily seen to be combinatorially valid; the partial map given by the pointed reduction can be made total by mapping unmapped vertices to those in $\llbracket i_{X}\rrbracket$. Then, by completeness we have $i_{X}\times t\leq u$ derivable, and since $1\leq_{\bullet}i_{X}$, we a fortiori have $t=1\times t\leq_{\bullet}i_{X}\times t\leq u$ derivable. ∎
5 Introducing joins and parallelization
Joins
{mdframed}
$\begin{array}[]{cr}a\leq a\sqcup b\qquad b\leq a\sqcup b&\text{$\sqcup$ is an upper bound}\\ b\leq a\;\wedge\;c\leq a~{}\Rightarrow~{}b\sqcup c\leq a&\text{$\sqcup$ is theleast ub}\\ a\sqcap(b\sqcup c)=a\sqcap b\sqcup a\sqcap c&\text{distributivity of $\sqcap$}\\ a\times(b\sqcup c)=a\times b\sqcup a\times c&\text{distributivity of $\times$}\\ \end{array}$
Parallelization {mdframed}
$\begin{array}[]{cr}a\leq a^{*}\qquad(a^{*})^{*}\leq a^{*}&\text{closure operator}\\ a^{*}\times a^{*}\leq a^{*}&\text{duplicable}\\ (a\sqcup b)^{*}=a^{*}\times b^{*}&\text{interaction with $\sqcup$}\\ (a\sqcap b)^{*}=a^{*}\sqcap b^{*}&\text{interaction with $\sqcap$}\\ (a\times b)^{*}=1\sqcup a\times a^{*}\times b\times b^{*}&\text{interaction with $\times$}\\ 1^{*}=1&\text{interaction with $1$}\\ \end{array}$
Figure 2: Axioms for $\sqcup$ and $(-)^{*}$
We extend our axiomatization of $(\mathfrak{W},\sqcap,\times,1)$ to an axiomatization of $(\mathfrak{W},\sqcap,\times,1,\sqcup,(-)^{*})$ as shown in Figure 2. This list of axioms is not free of redundancy. For example, duplicability of $(-)^{*}$ follows from $a=a\sqcup a$ and the interaction of $(-)^{*}$ with $\sqcup$. We will show that a complete axiomatization of $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$ yields a complete axiomatization of $(\mathfrak{W}_{\bullet},\sqcap,\times,1,\sqcup,(-)^{*})$.
Observe that the distributivity rules for $\sqcap$, $\times$, and $(-)^{*}$ in Figure 2 allow us to rewrite every term $t$ as a join $t=\bigsqcup_{i}t_{i}$ where the $t_{i}$s are $\sqcup$-free terms such that parallelization is only applied to variables. Distributing $\sqcap$ and $\times$ over $\sqcup$ can increase the size of the term exponentially.
However, we can re-write any term in linear time into one where parallelization is applied only to variables. This is obvious in the pointed case, since the rule $(a\times b)^{*}=1\sqcup a\times a^{*}\times b\times b^{*}$ together with pointedness yields $(a\times b)^{*}=a^{*}\times b^{*}$. In general, an easy induction shows that any term of the form $t=\left(\bigtimes_{i=1}^{n}t_{i}\right)^{*}$ is derivably equivalent to $1\sqcup\bigtimes_{i=1}^{n}\left(t_{i}\times t_{i}^{*}\right)$ via a linear-size derivation.
First, we extend the combinatorial criterion to terms with the additional connectives. With our axioms, we can without loss of generality assume that parallelizations are only applied to variables; terms that do not satisfy this can be rewritten to terms that do so in linear time to terms that are provably equivalent. We make this assumption below when defining the notion of combinatorial reduction for those terms.
[…]
For a term $t$ over $(\mathfrak{W}_{\bullet},\sqcap,\times,\sqcup,(-)^{*})$, we define the set $\mathrm{Slice}(t)$ of slices of $t$ by induction as follows:
  • If $t$ is $\sqcup$-free, then $\mathrm{Slice}(t)=\{t\}$.
  • $\mathrm{Slice}(t\sqcup u)=\mathrm{Slice}(t)\cup\mathrm{Slice}(u)$.
[…]
    $\mathrm{Slice}(t\otimes u)=\mathrm{Slice}(t)\otimes\mathrm{Slice}(u):=\left\{t^{\prime}\otimes u^{\prime}\mid t^{\prime}\in\mathrm{Slice}(t),u^{\prime}\in\mathrm{Slice}(u)\right\}.$
Any term $t^{\prime}\in\mathrm{Slice}(t)$ is $\sqcup$-free and can hence be associated with a graph with parallelization information $\llbracket t^{\prime}\rrbracket$.
Definition 19.
Let $t$ and $u$ be terms over $(\mathfrak{W}_{\bullet},\sqcap,\times,\sqcup,(-)^{*})$. A combinatorial reduction from $t$ to $u$ is given by a map $s\colon\mathrm{Slice}(t)\to\mathrm{Slice}(u)$ and a family of relations $f_{t^{\prime}}\subseteq V_{s(t^{\prime})}\times V_{t^{\prime}}$ indexed over the slices $t^{\prime}$ of $t$, such that the following conditions are satisfied:
  1. 1.
    For all $t^{\prime}\in\mathrm{Slice}(t)$, if $p_{s(t^{\prime})}(x)=0$ then $(x,y),(x,y^{\prime})\in f_{t^{\prime}}$ implies $y=y^{\prime}$ ($f_{t^{\prime}}$ is single-valued on non-parallelized variables).
  2. 2.
    For all $t^{\prime}\in\mathrm{Slice}(t)$, if $(x,y)\in f_{t^{\prime}}$ then $c_{t^{\prime}}(y)=c_{s(t^{\prime})}(x)$ ($f_{t^{\prime}}$ respects colours).
  3. 3.
    For all $t^{\prime}\in\mathrm{Slice}(t)$, if $(x,y)\in f_{t^{\prime}}$ then $p_{t^{\prime}}(y)\leq p_{s(t^{\prime})}(x)$ ($f_{t^{\prime}}$ respects parallelization information).
  4. 4.
    For all $t^{\prime}\in\mathrm{Slice}(t)$ and all maximal cliques $M$ in $\llbracket s(t^{\prime})\rrbracket$ the image of $M$ under $f_{t^{\prime}}$ contains a maximal clique of $t^{\prime}$.
We say that a pointed inequality $t\leq_{\bullet}u$ is combinatorially valid if there exists a corresponding combinatorial reduction between the two terms.
Lemma 20.
Combinatorial and universal validity coincide for terms over $(\mathfrak{W}_{\bullet},\sqcap,\sqcup,\times,(-)^{*})$.
Proof.
The structure of the proof is identical to that of the proof of Lemma 9.
We introduce generic validity of formulas just as we did for formulas over $(\mathfrak{W}_{\bullet},\sqcap,\times)$. It is clear that universal validity implies generic validity.
To show that that generic validity implies combinatorial validity one starts with a Weihrauch reduction $g_{t}\leq_{W}g_{u}$ where $g_{t}$ and $g_{u}$ are “generic interpretations” of terms $t$ and $u$ respectively. By a straightforward induction on the definitions of $\times$, $\sqcap$, $\sqcup$, and $\mathrm{Slice}$, one obtains that the forward Weihrauch reduction yields a combinatorial reduction from $t$ to $u$, whose correctness is witnessed by the backwards reduction. To prove the latter part, one queries in $g_{t}$ the generic Weihrauch degrees associated with different occurrences of non-parallelized variables for pairwise distinct integers. The parallelized degrees associated with occurrences of parallelized variables are queried for pairwise distinct integers that are also pairwise distinct from all the queries to non-parallelized variables. Further, one ensures that every parallelized degree associated with an occurrence of a parallelized variable is queried for a number of inputs that is strictly larger than the number of occurrences of variables in $u$.
Finally, it is fairly straightforward to construct a Weihrauch reduction from a combinatorial reduction, analogously to Lemma 9. ∎
Lemma 20 allows us to extend a complete axiomatization for $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$ to a complete axiomatization for $(\mathfrak{W}_{\bullet},\sqcap,\times,1,\sqcup,(-)^{*})$ by adding the axioms in Figure 2:
Proposition 21.
Let $\Gamma$ be a complete set of axioms for $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$. Let $\Delta$ denote the set of axioms from Figure 2. Then $\Gamma\cup\Delta$ is a complete set of axioms for $(\mathfrak{W}_{\bullet},\sqcap,\times,1,\sqcup,(-)^{*})$.
Proof.
We proceed in stages, starting from the complete axiomatization of $(\mathfrak{W}_{\bullet},\sqcap,\times)$ and improving to a complete axiomatization of $(\mathfrak{W}_{\bullet},\sqcap,\times,1,\sqcup,(-)^{*})$ by claiming that $\Gamma\cup\Delta$ is a complete axiomatization of the substructures obtained by adding the connectives $(-)^{*}$, $\sqcup$ and $1$ successively.
Adding parallelization:
Assume we have $t\leq_{\bullet}u$ universally valid in $(\mathfrak{W}_{\bullet},\sqcap,\times,(-)^{*})$. Without loss of generality, we can assume that $(-)^{*}$ is only applied to variables in $t$ and $u$. Then, by Lemma 20, there exists a combinatorial reduction from $t$ to $u$. Since both terms are $\sqcup$-free, this reduction is given by a single relation $f_{t}\subseteq V_{u}\times V_{t}$ as in Definition 19. Consider those vertices $x\in V_{u}$ that are related by $f_{t}$ to a set of vertices $\{y_{1},\dots,y_{m}\}$ in $t$ with $m>1$. Such vertices necessarily correspond to subterms of $u$ of the form $a^{*}$, where $a$ is a variable. Construct a new graph $G^{\prime}$ from $\llbracket u\rrbracket$ by replacing all such vertices $x$ in $\llbracket u\rrbracket$ by a clique of $m$ vertices $x_{1},\dots,x_{m}$, where $m$ is the size of the image of $\{x\}$ under $f_{t}$, all coloured with the same colour as $x$ and with $p(x_{i})=1$ for all $i$. We have $G^{\prime}=\llbracket u^{\prime}\rrbracket$, where $u^{\prime}$ is obtained from $u$ by replacing certain subterms of the form $a^{*}$ with subterms of the form $a^{*}\times\dots\times a^{*}$. The equality $u=_{\bullet}u^{\prime}$ is derivable in $\Gamma\cup\Delta$ by employing relevance of $\times$ and the duplication axiom from Figure 2. By construction, the combinatorial reduction $f_{t}$ from $t$ to $u$ can be made into a single-valued combinatorial reduction $g$ from $t$ to $u^{\prime}$. Fix a mapping $P\colon\mathcal{V}\to\mathcal{V}$ between variables, sending the variables that occur in $u$ or $t$ to fresh distinct variables, not occurring in either term. Let $\hat{t}$ denote the term which is obtained from $t$ by replacing all subterms of $t$ of the form $a^{*}$ for a variable $a$ with the variable $P(a)$. Let $\hat{u}$ denote the term which is obtained from $u^{\prime}$ by replacing all subterms of $u^{\prime}$ of the form $a^{*}$ for a variable $a$ as follows: If the vertex in $\llbracket u^{\prime}\rrbracket$ corresponding to the subterm gets mapped by $g$ to a vertex which corresponds to a term of the form $a^{*}$, replace the subterm with the variable $P(a)$. If the vertex $\llbracket u^{\prime}\rrbracket$ corresponding to the subterm gets mapped by $g$ to a vertex which corresponds to a term of the form $a$, replace the term with the variable $a$. By construction, the terms $\hat{u}$ and $\hat{t}$ belong to the language of the fragment $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$, and we have a combinatorial reduction from $\hat{t}$ to $\hat{u}$. By assumption, this yields a derivation over $\Gamma$ of the inequality $\hat{t}\leq_{\bullet}\hat{u}$. By reversing the substitution we have performed everywhere in the derivation, we obtain a derivation of the inequality $t\leq_{\bullet}u^{\prime}$. Since $u^{\prime}=_{\bullet}u$ is derivable, so is $t\leq_{\bullet}u$ as required.
Adding joins:
Now, consider the full fragment $(\mathfrak{W}_{\bullet},\sqcap,\times,\sqcup,(-)^{*})$. Suppose that an inequality $t\leq_{\bullet}u$ in this fragment is universally valid. Then by the above considerations we obtain a derivably equivalent formula of the form $\bigsqcup_{i}t_{i}\leq_{\bullet}\bigsqcup_{j}u_{j}$ where the $t_{i}$s and $u_{j}$s are $\sqcup$-free and finite parallelizations are applied only to variables. By the axiom stating that $\sqcup$ is the least upper bound, this further reduces to deriving the universally valid formula $t_{i}\leq_{\bullet}\bigsqcup_{j}u_{j}$ for all $i$.
Now, observe that $t\leq_{\bullet}u\sqcup v$ is universally valid for a $\sqcup$-free term $t$ as above if and only if $t\leq_{\bullet}u$ or $t\leq_{\bullet}v$ is universally valid. Indeed, by Lemma 20, universal validity of an inequality is equivalent to the existence of a combinatorial reduction. By definition, a combinatorial reduction from $t$ to $u\sqcup v$ for $\sqcup$-free $t$ immediately yields a combinatorial reduction from $t$ to $u$ or from $t$ to $v$, since we have $\mathrm{Slice}(t)=\{t\}$ and $\mathrm{Slice}(u\sqcup v)=\mathrm{Slice}(u)\cup\mathrm{Slice}(v)$
Hence, from the universal validity of $t_{i}\leq_{\bullet}\bigsqcup_{j}u_{j}$ we obtain the universal validity of $t_{i}\leq_{\bullet}u_{j}$ for some $j$, which is over $(\mathfrak{W}_{\bullet},\sqcap,\times,(-)^{*})$ and thus derivable from $\Gamma\cup\Delta$.
Adding $1$:
Given a term $t$ over the full signature, it is easy to check whether it is equivalent to $1$ and then prove from $\Gamma\cup\Delta$ that it is equivalent to it if need be, or to derive $t=_{\bullet}t^{\prime}$ with $t^{\prime}$ free of $1$. Then if we are to prove a universally valid inequality $t\leq_{\bullet}u$ we have three cases:
  • both $t$ and $u$ are equivalent to $1$-free terms hence we derive $t=_{\bullet}t^{\prime}\leq_{\bullet}u^{\prime}=_{\bullet}u$ using our previous point
  • if $t=_{\bullet}1$, then we can then derive $t=_{\bullet}1\leq_{\bullet}u$ where the latter is derived by recurring over $u$
  • if $u=_{\bullet}1$, then it is also the case that $t=_{\bullet}1$ since $1$ is the bottom element of $\mathfrak{W}_{\bullet}$, and we can derive $u=_{\bullet}1=_{\bullet}t$.
Finally, we observe that we can extend our notion of combinatorial validity to non-pointed terms over $(\mathfrak{W},\sqcap,\times,(-)^{*},1)$ by requiring that the relations $f_{t^{\prime}}$ be left-total and handling $1$ as indicated at the end of Section 3. Using this, we obtain an analogous result to Proposition 21 for the non-pointed case. The proof is almost identical, so we will not spell it out in detail.
6 Complexity of deciding validity
We turn to the problem of determining the validity of an inequality over $(\mathfrak{W}_{\bullet},\sqcap,\times,1,(-)^{*})$ or $(\mathfrak{W}_{\bullet},\sqcap,\times,1,\sqcup,(-)^{*})$. We first consider the $\sqcup$-free case:
Theorem 22.
The problem “is the inequality $t\leq_{\bullet}u$ valid?” in the structure $(\mathfrak{W}_{\bullet},\sqcap,\times)$ is $\Sigma^{p}_{2}$-complete.
Proof.
In view of Lemma 20 we can consider the equivalent problem of deciding the existence of a combinatorial reduction from $t$ to $u$, i.e. a $\bullet$-reduction from $\llbracket t\rrbracket$ to $\llbracket u\rrbracket$.
For containment in $\Sigma_{2}^{p}$, note that the existence of a reduction can be essentially written as
$\exists f:V_{u}\rightharpoonup V_{t}.~{}~{}\underbrace{\forall~{}C\text{ max. clique of }\llbracket u\rrbracket.~{}~{}\overbrace{f(C)\text{ contains a max. clique of }\llbracket u\rrbracket}^{\text{polytime}}}_{\text{coNP}}.$
In the above, all quantifiers involve polynomial-sized objects relative to $G$ and $H$. It hence suffices to count quantifiers and argue that we may indeed check the inner condition in polynomial time.
In general, checking whether a set of vertices contains a maximal clique may be hard. However, in our case, since we are working with denotations of terms, we are not working with general graphs, but cographs. In this case, there is a straightforward algorithm for deciding this problem. For a term $t$, let the predicate $\operatorname{MClique}_{t}(s,V)$ for subterms $s$ of $t$ and sets $V$ of occurrences of variables in $t$ be defined as follows:
  1. 1.
    For an occurrence $v$ of a variable, $\operatorname{MClique}_{t}(v,V)$ holds true if and only if $v\in V$.
  2. 2.
    $\operatorname{MClique}_{t}(s\sqcap u,V)=\operatorname{MClique}_{t}(s,V)\lor\operatorname{MClique}_{t}(u,V)$.
  3. 3.
    $\operatorname{MClique}_{t}(s\times u,V)=\operatorname{MClique}_{t}(s,V)\land\operatorname{MClique}_{t}(u,V)$.
If we represent the set $V$ appropriately, the predicate $\operatorname{MClique}_{t}(s,V)$ can be evaluated in essentially linear time. It is clear that $\operatorname{MClique}_{t}(s,V)$ holds true if and only if the set $V$ contains a maximal clique in $\llbracket t\rrbracket$.
To show $\Sigma_{2}^{p}$-hardness, we reduce in polynomial time TQBF restricted to $\Sigma_{2}$ formulas to our problem. Given a quantified boolean formula
$\exists x_{1},\ldots,x_{k}\forall y_{1},\ldots,y_{n}~{}\bigwedge_{i\in I}\bigvee_{j\in J_{i}}\ell_{j},$
with the $\ell_{j}$s being literals generated from $\{x_{1},\ldots,x_{k},y_{1},\ldots,y_{n}\}$, define $K=1+\sum\limits_{i\in I}|J_{i}|$. We may then produce in polynomial time the following terms whose variables are literals (where powers are to be understood as iterated $\times$):
  • $E_{L}=\bigtimes\limits_{m=1}^{k}x_{m}^{K}\sqcap\overline{x_{m}}^{K}$ and $E_{R}=\bigtimes\limits_{m=1}^{k}x_{m}^{K}\times\overline{x_{m}}^{K}$
  • $A=\bigtimes\limits_{m=1}^{n}y_{m}^{K}\sqcap\overline{y_{m}}^{K}$
  • $F=\bigtimes\limits_{i\in I}\bigsqcap\limits_{j\in J_{i}}\ell_{j}$
To conclude, we only need to argue that the formula is satisfiable if and only if the inequality
$E_{L}\times F\leq_{\bullet}E_{R}\times A$
is valid.
Assume that the formula is satisfiable. Then there exists a valuation $\rho\colon\{x_{1},\dots,x_{k}\}\to\{0,1\}$ of the $x_{i}$s such that for all valuations $\theta\colon\{y_{1},\dots,y_{n}\}\to\{0,1\}$ of the $y_{j}$s, the combined valuation $(\rho,\theta)$ renders the formula true.
Pick any $\rho$ as above. Then there exists a partial mapping $f\colon E_{R}\times A\rightharpoonup E_{L}\times F$ with the following properties:
  1. 1.
    If $\rho(x_{i})=1$, the map $f$ sends all occurrences of $\overline{x_{i}}$ in $E_{R}$ to an occurrence of $\overline{x_{i}}$ in $E_{L}$, such that the clique $\overline{x_{i}}^{K}$ in $E_{L}$ is contained in the image of $f$ under he clique $\overline{x_{i}}^{K}$ in $E_{R}$. Further, the image of the clique $x_{i}^{K}$ in $E_{R}$ covers all occurrences of $x_{i}$ in $F$.
  2. 2.
    If $\rho(x_{i})=0$, the map $f$ has the same properties, with the roles of $x_{i}$ and $\overline{x_{i}}$ reversed.
  3. 3.
    The image of the clique $y_{m}^{K}$ covers all occurrences of $y_{m}$ in $F$.
  4. 4.
    The image of the clique $\overline{y_{m}}^{k}$ covers all occurrences of $\overline{y_{m}}$ in $F$.
The mapping $f$ exists essentially because $K$ is chosen large enough to ensure that we can cover all occurrences of variables in $F$ as described above.
Now, we claim that $f$ is a combinatorial reduction from $E_{L}\times F$ to $E_{R}\times A$. Let $C$ be maximal clique in $E_{R}\times A$. Then $C$ is of the form $C=E_{R}\times C^{\prime}$, where $C^{\prime}$ is a maximal clique in $A$. Further, $C^{\prime}$ is of the form
$C^{\prime}=\bigtimes\limits_{m=1}^{n}z_{m}^{K}$
where $z_{m}\in\{y_{m},\overline{y_{m}}\}$. Let $\theta\colon\{y_{1},\dots,y_{n}\}\to\{0,1\}$ be the valuation that sends $y_{j}$ to $1$ if $z_{j}=y_{j}$ and to 00 if $z_{j}=\overline{y_{j}}$. Then, by assumption, some conjunction of literals $\bigwedge_{i\in I}\ell_{j(i)}$ holds true under the valuations $\rho$ and $\theta$. Consider the corresponding maximal clique ×$\bigtimes_{i\in I}\ell_{j(i)}$ in $F$. Let $E_{L}^{\prime}$ be the maximal clique ×$\bigtimes_{m=1}^{k}w_{m}^{K}$ of $E_{L}$ where $w_{m}\in\{x_{m},\overline{x_{m}}\}$ and $w_{m}=x_{m}$ if and only if $\rho(x_{m})=0$. Then $E_{L}^{\prime}\times\bigtimes_{i\in I}\ell_{j(i)}$ is a maximal clique in $E_{L}\times F$. By construction, every vertex in $E_{L}$ is the image of a vertex in $E_{R}$ under $f$.
Now, by assumption, $\ell_{j(i)}$ is consistent with $\rho$ and $\theta$ in the sense that if $\ell_{j(i)}\in\{x_{m},\overline{x_{m}}\}$, then $\ell_{j}(i)=x_{m}$ if and only if $\rho(x_{m})=1$ and similarly for $\theta$ and $\ell_{j(i)}\in\{y_{m},\overline{y_{m}}\}$. But this implies by construction that every $\ell_{j(i)}\in\{y_{m},\overline{y}_{m}\}$ is the image of a vertex in $C^{\prime}$ under $f$ and every $\ell_{j(i)}\in\{x_{m},\overline{x}_{m}\}$ is the image of a vertex in $E_{R}$ under $f$. This shows that $f$ is a combinatorial reduction.
Conversely, assume that a combinatorial reduction exists between $E_{L}\times F$ and $E_{R}\times A$, witnessed by a partial map
$f\colon\llbracket E_{R}\times A\rrbracket\rightharpoonup\llbracket E_{L}\times F\rrbracket.$
Observe that for all $m\in\{1,\dots,k\}$, either $f$ maps the clique $x_{m}^{K}$ of $\llbracket E_{R}\rrbracket$ into $\llbracket E_{L}\rrbracket$, or it maps the clique $\overline{x_{m}}^{K}$ of $\llbracket E_{R}\rrbracket$ into $\llbracket E_{L}\rrbracket$.
Hence, we can define a valuation $\rho\colon\{x_{1},\dots,x_{k}\}\to\{0,1\}$ by letting $\rho(x_{i})=1$ if and only if the clique $\overline{x_{m}}^{K}$ of $\llbracket E_{R}\rrbracket$ is mapped into $\llbracket E_{L}\rrbracket$. It is now quite straightforward to check that $\rho$ is a satisfying assignment for the formula. ∎
Our argument for containment in $\Sigma_{2}^{p}$ can easily be adapted to formulas with parallelization applied to variables. Using the observation that any formula with parallelization can be re-written in this shape in linear time, we obtain:
Corollary 23.
The problem “is the inequality $t\leq_{\bullet}u$ valid?” in the structure $(\mathfrak{W}_{\bullet},\sqcap,\times,(-)^{*})$ is $\Sigma^{p}_{2}$-complete.
The proof of Theorem 22 (together with the observation underlying Corollary 23) yields that validity of inequalities in $(\mathfrak{W}_{\bullet},\sqcap,\times,1,\sqcup,(-)^{*})$ is contained in $\Pi_{3}^{p}$ — the additional universal quantifier being used to quantify over the slices. This additional quantifier alternation cannot be avoided:
Theorem 24.
The problem “is the inequality $t\leq_{\bullet}u$ valid?” in the structure $(\mathfrak{W}_{\bullet},\sqcap,\times,1,\sqcup)$ is $\Pi^{p}_{3}$-complete.
Proof idea.
Inclusion in $\Pi^{p}_{3}$ is straightforward, following the proof of Theorem 22. For the hardness part, we employ a similar construction to that in the proof of Theorem 22. Given a quantified boolean formula
$\forall z_{1},\ldots,z_{l}\exists x_{1},\ldots,x_{k}\forall y_{1},\ldots,y_{n}~{}\bigwedge_{i\in I}\bigvee_{j\in J_{i}}l_{j}$
we construct the inequality
$Z_{L}\times E_{L}\times F\leq_{\bullet}Z_{R}\times E_{R}\times A$
where $E_{L}$, $F$, $E_{R}$ and $A$ are taken as in the proof of Theorem 22 ($F$ might contain literals $z_{i}$ and $\overline{z_{i}}$), and where
$Z_{L}=\bigtimes_{i=1}^{l}(z_{i}\sqcup\overline{z_{i}})\qquad\text{and}\qquad Z_{R}=\bigtimes_{i=1}^{l}(z_{i}^{K}\sqcup\overline{z_{i}}^{K}).$
7 Towards a proof of completeness
We do not know whether our axiom system yields a complete axiomatization of $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$. As a first possible step towards a potential completeness proof, we show that we can restrict our attention to inequalities $t\leq_{\bullet}u$ where $t$ and $u$ each have pairwise distinct variables and the combinatorial reduction witness is a total bijective map.
[…]
Proposition 25.
Let $t$ and $u$ be terms over $(\mathfrak{W}_{\bullet},\sqcap,\times,1)$. Assume that $t\leq_{\bullet}u$. Then there exist terms $t^{\prime}$ and $u^{\prime}$ such that $t\leq_{\bullet}t^{\prime}\leq_{\bullet}u^{\prime}\leq_{\bullet}u$, the inequalities $t\leq_{\bullet}t^{\prime}$ and $u^{\prime}\leq_{\bullet}u$ are derivable from the axioms in Figure 1, and such that there exists a (total) bijective combinatorial reduction from $t^{\prime}$ to $u^{\prime}$.
Proof.
As argued at the end of Section 3, we can assume without limitation of generality that $t$ and $u$ are either $1$ or do not contain a $1$. If either term is $1$, the claim is trivial. Otherwise, there exists a combinatorial reduction $f\colon V_{u}\rightharpoonup V_{t}$.
Consider a vertex $v$ in $\llbracket t\rrbracket$ not in the range of $f$. Consider the path $\langle o_{1},\dots,o_{m},v\rangle$ in the (tree associated with the) term $t$ leading to $v$. If $o_{1}=\dots=o_{m}=\times$, then every maximal clique in $\llbracket t\rrbracket$ contains $v$, contradicting the assumption that $f$ is a combinatorial reduction. Hence, there exists a maximal $j\in\{1,\dots,m\}$ with $o_{j}=\sqcap$. Let $t^{\prime}$ be the term which is obtained from $t$ by removing the sub-tree with root $o_{j+1}$ (letting $o_{m+1}=v$). Then we can derive $t\leq_{\bullet}t^{\prime}$ since meets are lower bounds. The maximal cliques of $\llbracket t^{\prime}\rrbracket$ are the maximal cliques of $t$ that do not contain $v$. Since no maximal clique that contains $v$ contributes to $f$ being a combinatorial reduction, the co-restriction of $f$ to $V_{t^{\prime}}$ remains a combinatorial reduction. We can apply the above construction to remove all vertices outside the range of $f$ to ensure that $f$ is a surjective partial map – potentially with smaller domain.
If $f$ is not injective, say $f^{-1}(\{v\})=\{w_{1},\dots,w_{m}\}$, we can replace $v$ in $\llbracket t\rrbracket$ with the $m$-fold meet $v\sqcap\dots\sqcap v$ (coloured with the same colour as $v$) and send each $w_{i}$ to a separate occurrence of $v$. It is straightforward to check that this yields a combinatorial reduction from $u$ to a term $t^{\prime}$ which is derivably equivalent to $t$.
In the case where $f$ is not total with $v\in V_{u}\setminus\operatorname{dom}(f)$, we consider the path $\langle o_{1},\dots,o_{m},v\rangle$ in the (tree associated with the) term $u$ leading to $v$. There must exist a maximal $j\in\{1,\dots,m\}$ with $o_{j}=\times$. Removing in $u$ the the sub-tree with root $o_{j+1}$ (letting $o_{m+1}=v$) yields a derivably smaller term $w$ such that $f$ is still a combinatorial reduction from $\llbracket w\rrbracket$ to $\llbracket t\rrbracket$. Applying this construction to all vertices outside $\operatorname{dom}(f)$, we obtain a term $u^{\prime}$ as claimed. ∎
Proposition 25 implies that a complete axiomatization for formulas of the form $t\leq_{\bullet}u$ where $t$ and $u$ both have pairwise distinct variables yields a complete axiomatization of the whole theory: if $t$ and $u$ are arbitrary terms such that $t\leq_{\bullet}u$ is universally valid, there exists without loss of generality a bijective combinatorial reduction from $u$ to $t$. This allows us to transport an injective colouring of the vertices of $\llbracket u\rrbracket$ to an injective colouring of the vertices of $\llbracket t\rrbracket$, yielding a valid and hence derivable inequality $t^{\prime}\leq_{\bullet}u^{\prime}$ between terms with pairwise distinct variables. Substituting back the original variable names in the derivation of $t^{\prime}\leq_{\bullet}u^{\prime}$ yields a derivation of $t\leq_{\bullet}u$.
Proposition 26.
Our axiom system is complete for inequalities $t\leq_{\bullet}u$ where $u$ is square-free.
Proof.
By the remark following Proposition 25 we may assume that the variables are pairwise distinct in both $t$ and $u$. We show the claim by induction over $u$. If $u$ is a variable, the constant $1$, or of the form $u_{0}\sqcap u_{1}$, this is obvious.
The remaining case is $u=u^{\prime}\times x$. An easy induction over $t$ shows that we can derive $t\leq_{\bullet}t[1/x]\times x$. Since $x$ does not appear in $u^{\prime}$, we must have $t[1/x]\leq_{\bullet}u^{\prime}$. By the induction hypothesis, this last inequality is derivable from our axioms. Now, combining this derivation with monotonicity of $\times$ yields a derivation of $t[1/x]\times x\leq_{\bullet}u^{\prime}\times u$ which in total yields a derivation of $t\leq_{\bullet}u$. ∎
Acknowledgments
We are grateful to Manlio Valenti and Lê Thành Dũng Nguyễn for discussions, and also to Matteo Acclavio for some pointers to the literature.